let f be the function defined by f(x)=4x-3. Find f(4), f(1/4), f(0), f(a), and f(a+1)
Found 2 solutions by solver91311, mastergerwe:
Answer by solver91311(24713)











Always do any indicated arithmetic if possible.


John

My calculator said it, I believe it, that settles it


Answer by mastergerwe(4)

f(4)=4(4)-3 = 13
f(1/4)=4(1/4)-3=-2
f(0)=4(0)-3=-3
f(a)=4(a)-3=4a-3
f(a+1)=4(a+1)-3=4a-1
:)